# Double Landmines: Invisible Textual Backdoor Attacks based on Dual-Trigger

### v1

@@ V-B1 Robustness against ONION
OLID Benign - 63.21(-0.82) - 51.69(+0.35) - 44.82(-12.46)
BadNet 38.18(-61.59) 81.72(-3.5) 45.63(-54.25) 78.23(-3.38) 35.16(-64.72) 80.68(-2.91)
@@
Syntactic 95.16(-4.68) 78.93(-5.47) 83.87(-12.42) 79.51(-5.36) 95.32(-4.2) 75.44(-4.89)
Dual-Trigger 99.07(-0.93) 72.88(-10.71) 96.97(-2.68) 67.05(-13.39) 93.02(-3.49) 75.67(-8.26)
AG’s News Benign - 85.64(-0.48) - 51.64(-0.65) - 56.32(-0.25)
BadNet 47.42(-52.58) 79.57(-15.79) 32.46(-67.54) 91.96(-2.07) 64.63(-35.37) 64.56(-5.5)
@@
Dual-Trigger 95.96(-3.74) 89.91(-5.66) 98.83(-1.03) 80.16(-13.81) 94.63(-5.12) 70.06(-6.57)
TABLE VIII: Impact of two sentence-level defenses on backdoor attack performance of different backdoor attack methods (results are based on the SST-2 dataset).
@@
Back-translation Paraphrasing Benign - 92.09(-3.3) - 89.73(+0.22) - 84.51(+4.22)
BadNet 56.67(43.33) 94.29(-3.35) 54.42(-45.53) 93.96(-3.13) 56.56(-43.44) 92.86(-2.8)
@@
Amphibia 91.65(-8.02) 71.55(-25.98) 72.98(-26.2) 91.59(-4.51) 79.02(-20.76) 87.37(-8.51)
Syntactic Structure Alteration Benign - 93.03(-2.36) - 87.10(-2.41) - 77.92(-2.37)
@@
Amphibia 86.55(-13.12) 89.57(-7.96) 77.10(-22.08) 93.79(-2.31) 82.76(-17.02) 91.98(-3.9)
V-B2 Robustness against Sentence-Level Defenses
@@
Table VIII shows the performance of different backdoor attacks on the SST-2 dataset after adopting two sentence-level defense strategies. As can be seen from the table, although the first defense strategy based on Back-translation Paraphrasing can effectively mitigate the impact of the three baseline attacks, it is still insufficient to resist the Dual-Trigger attack method, and our method still achieves the best ASR. In contrast, the second defense measure focuses on eliminating syntactic triggers and successfully achieves a relatively ideal defense level. Even so, it will cause a similar or even greater drop in the ASR of the baseline. The above findings fully demonstrate the strong robustness of Dual-Trigger against sentence-level defense.
VI Poisoned Sample Example
In Table IX, we list some Dual-Trigger poisoned samples designed based on syntactic structure and subjunctive mood and their corresponding original normal samples. By observing these poisoned samples, we can find that they are not only fluent and natural, but also highly concealed, which makes it difficult to identify them by both automated tools and manual review.
TABLE IX: Examples of poisoned samples embedded with Dual-Trigger and the corresponding original normal samples.
@@ VI Poisoned Sample Example
VII CONCLUSION AND FUTURE WORK

### v2

@@ 5.2.1 Robustness against ONION
OLID Benign - 79.28(-0.82) - 65.89(-2.45) - 55.76(-1.52)
BadNet 38.18(-61.59) 81.72(-3.50) 45.63(-54.25) 78.23(-3.38) 35.16(-64.72) 80.68(-2.91)
@@
Syntactic 95.16(-4.68) 78.93(-5.47) 83.87(-12.42) 79.51(-5.36) 95.32(-4.20) 75.44(-4.89)
Dual-Trigger 99.07(-0.93) 79.05(-4.54) 96.97(-2.68) 77.88(-2.56) 93.02(-3.49) 75.67(-8.26)
AG’s News Benign - 85.64(-0.48) - 73.96(-0.12) - 56.32(-0.25)
BadNet 47.42(-52.58) 91.38(-3.98) 32.46(-67.54) 91.96(-2.07) 64.63(-35.37) 91.51(-3.22)
@@
Dual-Trigger 95.96(-3.74) 89.91(-5.66) 98.83(-1.03) 87.19(-6.78) 94.63(-5.37) 90.20(-3.76)
Table 8: Impact of two sentence-level defenses on backdoor attack performance of different backdoor attack methods (results are based on the SST-2 dataset).
@@
Back-translation Paraphrasing Benign - 92.09(-3.30) - 90.61(-1.65) - 89.90(-0.54)
BadNet 56.67(43.33) 94.29(-3.35) 54.42(-45.53) 93.96(-3.13) 56.56(-43.44) 92.86(-2.80)
@@
Amphibia 91.65(-8.02) 90.88(-6.65) 72.98(-26.20) 91.59(-4.51) 79.02(-20.76) 87.37(-8.51)
Syntactic Structure Alteration Benign - 93.03(-2.36) - 90.23(-2.03) - 88.36(-2.08)
@@
Amphibia 86.55(-13.12) 93.36(-4.17) 77.10(-22.08) 93.79(-2.31) 82.76(-17.02) 91.98(-3.90)
5.2.2 Robustness against Sentence-Level Defenses
@@
Table 8 shows the performance of different backdoor attacks on the SST-2 dataset after adopting two sentence-level defense strategies. As can be seen from the table, although the first defense strategy based on Back-translation Paraphrasing can effectively mitigate the impact of the three baseline attacks, it is still insufficient to resist the dual-trigger attack method, and our method still achieves the best ASR. In contrast, the second defense measure focuses on eliminating syntactic triggers and successfully achieves a relatively ideal defense level. Even so, it will cause a similar or even greater drop in the ASR of the baseline. The above findings fully demonstrate the strong robustness of dual-trigger against sentence-level defense.
6 Poisoned Sample Example
In Table 9, we list some dual-trigger poisoned samples designed based on syntactic structure and subjunctive mood and their corresponding original normal samples. By observing these poisoned samples, we can find that they are not only fluent and natural, but also highly concealed, which makes it difficult to identify them by both automated tools and manual review.
Table 9: Examples of poisoned samples embedded with Dual-Trigger and the corresponding original normal samples.
@@ 6 Poisoned Sample Example
7 CONCLUSION AND FUTURE WORK